OpenStudy (anonymous):

Solve the following quadratic trig equations algebraically (without graphing technology) for the domain 0 ≤ x ≤ 360°. Use exact solutions whenever possible. 1. 6cos^2(θ) + cos(θ) - 1 = 0 2. 2 - 2sin^2(θ) = cos(θ) Thanks for your help

OpenStudy (tkhunny):

What's preventing you from solving these? For the first, simply factor. \((2\cos(\theta) + 1)(3\cos(\theta) - 1) = 0\). You could have done that! For the second, use the Pythagorean identity to convert the sine to cosine and it will be very similar to the first.